(i) How many ways can you save the files into the 6 folders if the temporal order in which

you save the individual files does not matter.

(ii) How many ways can you save the files into the 6 folders if the temporal order in which

you save the individual files matters.



1
Expert's answer
2021-12-08T14:29:05-0500

i)

"N=\begin{pmatrix}
 n+ m-1\\
 m
\end{pmatrix}"

where m is number of files,

n is number of folders

"N=\begin{pmatrix}
 6+ m-1\\
 m
\end{pmatrix}=\begin{pmatrix}
 5+ m\\
 m
\end{pmatrix}=\frac{(5+m)!}{5!m!}"


ii)

"N=n^m=6^m"